A triangular piece of glass is being cut so that the height of the triangle is 4 inches shorter than twice the base. If the area of the triangle is 80 square inches, how long is the height of the triangle?
Answer by checkley77(12844)

Area=bh/2
80=b(2b-4)/2
80=(2b^2-4b)/2 cross multiply
2b^2-4b=2*80
2b^2-4b=160
2b^2-4b-160=0
2(b^2-2b-80)=0
2(b-10)(b+8)=0
b-10=0
b=10 in.
h=2*10-4
h=20-4
h=16 in.
Proof:
80=10(2*10-4)/2
80=10(20-4)/2
80=10*16/2
80=160/2
80=80